If the temperature of the helium balloon were increased from 30°C to 35°C and the volume of the balloon only expanded from 0.47L to 1.55L, how much pressure would be inside the balloon?
1
Expert's answer
2020-04-22T09:19:44-0400

According to the gas law equation:

p1V1/T1 = p2V2/T2

where p1 - initial pressure, V1 - initial volume, T1 - initial temperature, p2 - final pressure, V2 - final volume, T2 - final temperature.

From here:

p1/p2 = V2T1/V1T2

As 30°C = 303.15 K and 35°C = 308.15 K:

p1/p2 = 1.55 L × 303.15 K / 0.47 L × 308.15 K = 3.24

As a result, the pressure in the ballon is 3.24 times lower compared to the initial pressure.